Need help with this equation please!
0.3x - 0.2 (3-5x) = -0.5 (x-6)
I tried solving and got 4 but I have seen where they change the equation like this
0.3x + -0.2(3+-5x)=-0.5(x+-6) I don't understand how you can put + then the neg? please explain
Thank you

Answer by ewatrrr(24785)

Multiply thru by 10 to rid the decimals...:)
3x - 2 (3-5x) = -5 (x-6)
I'm 'old fashion( 2negatives = plus)
3x- 6 + 10x = -5x + 30
18x = 36
x = 36/18 = 2
And...checking
3x - 2 (3-5x) = -5 (x-6)
6 -2(-7) = 6+ 14 = 20
....
-5 is the same as +(-5)
However: sooner or later will deal with 2 negatives = positive